OpenStudy (anonymous):

Write the equation of the circle in standard form.  Find the center, radius, intercepts, and graph the circle. x^2+y^2-8x-6y+9=0

OpenStudy (anonymous):

Complete the square on "x" and "y" together: (x^2 - 8x + 16) + (y^2 - 6y + 9) + 9 - 16 - 9 = 0 (x - 4)^2 + (y - 3)^2 = 4^2 So, the center is (4, 3). The radius is 4. To get the intercepts, you set each variable to "0", first "x" and then "y". (0 - 4)^2 + (y - 3)^2 = 4^2 -> y = 3 so, (0, 3) (x - 4)^2 + (0 - 3)^2 = 4^2 -> (x - 4)^2 = 7 x - 4 = +- sqrt(7) -> x = 4 +- sqrt(7)

OpenStudy (anonymous):

is the end of this x-4=+-^7

OpenStudy (anonymous):

That last intercept is : (4 +- sqrt[7], 0) that's 2 points because of the +-

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

Is there anything you want me to go over in more depth? @artishaalston

OpenStudy (anonymous):

this whole class lol

OpenStudy (anonymous):

: - )

OpenStudy (anonymous):

Here's a drawing that might help.

OpenStudy (anonymous):

Here, you can see that the y-axis is touched at only one point but the x-axis has those 2 points for intercepts.

OpenStudy (anonymous):

The key here was "completing the square" on x and y both at the same time. From there, it goes less rocky.

OpenStudy (anonymous):

Your general form for a circle will be: (x - h)^2 + (y - k)^2 = r^2 where (h, k) will be your center and "r" is your radius.

OpenStudy (anonymous):

Regarding your comment about "the whole class" which is good humor, will be helped by just showing up at Openstudy a lot. There are a lot of helpful people here, so no worries.

OpenStudy (anonymous):

If you need a more complete diagram:

OpenStudy (anonymous):

This particular problem, answer, and diagram might be helpful for questions about circles for the future.

OpenStudy (anonymous):

Anything else about this problem now?

OpenStudy (anonymous):

@artishaalston ?

OpenStudy (anonymous):

Good luck to you in all of your studies and so long for now. @artishaalston